OpenStudy (anonymous):

what is the number of atoms in 5.40 g Boron

OpenStudy (anoop27):

Find Moles First No. of Moles =5.40/10.81=1/2 use Mole to find out No. of Atoms No. of Atoms =1/2* Avogardo No.